OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. (1 point) y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Point slope form of line is given by: \[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

yes

OpenStudy (sleepyjess):

\(y-y_1=m(x-x_1)\) they give you the point (-1, 6) so you will use that for \(x_1,~y_1\)

OpenStudy (anonymous):

\((x_!, y_1)\) is the given point ie \((-1,6)\) and Slope \(m\) is \(3\).. Just plug in the values.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Just try it..

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

Excellent.. :)

OpenStudy (anonymous):

* \(m = -3\), I have written wrong there.. :(

OpenStudy (sleepyjess):

\(\Huge\frak\color{magenta}{B~}\Huge\frak\color{cyan}{is~}\Huge\frak\color{lime}{correct!}\)

OpenStudy (anonymous):

im right!?

OpenStudy (anonymous):

Haha thanks can i ask another?

OpenStudy (sleepyjess):

\(\Huge\cal\color{blue}{Yes...~you~are~correct!}\)

OpenStudy (anonymous):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 60 dumpster loads of waste removed. In week 10, there were 40 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. (2 points) f(x) = - 4x + 60 f(x) = 4x + 60 f(x) = - 4x + 80 f(x) = 4x + 80

OpenStudy (sleepyjess):

sure \(\ddot\smile\)

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0). (2 points) y - 4 = - (x + 6) y - 4 = 2(x + 6) y + 6 = - (x - 4) y + 6 = 2(x - 4) @sleepyjess

OpenStudy (anonymous):

For third one: First find lope by using : \[Slope (m) = \frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (sleepyjess):

first we need to find the slope do you know how to do that? @shmual

OpenStudy (anonymous):

Can you find Slope?